OpenStudy (cicilybailey14):

1 to 4 = y to 8 Medal :D Does Y= 4??

OpenStudy (texaschic101):

1 to 4 can be written as 1/4...same thing with y to 8....y/8 so we have a proportion : 1/4 = y/8 cross multiply 4y = 8 -- divide both sides by 4 y = 2 check.. 1/4 = 2/8 cross multiply 4 * 2 = 1 * 8 8 = 8 (correct) so y = 2